OpenStudy (anonymous):

A car(500kg) is being towed by a trailer(2000kg) with an acceleration of 2.4m/s^2. plz find the force applied to the road by car.

OpenStudy (anonymous):

To solve this you should try to make a free body diagram to show how gravity, normal force, the force on the car by the trailer, and the force of friction would interact. What force would then be applied to the road by the car?

OpenStudy (anonymous):

\[F =m \times a\] F=(500kg)*(2.4m/s^2)=1200N

OpenStudy (anonymous):

You need to know what type of force you are after and also what direction it is acting (vertical - horizontal)

OpenStudy (prajwal):

I think we should assume the surface is frictionless for no information is provided about that. Now it is obvious that the acceleration of car is equal to acceleration of the trailer. Mass of car= 500 kg Acceleration of car = 2.4 m/s^2 Force = ma = 500*2.4 = 1200 N

OpenStudy (anonymous):

I think that if we are assuming that there is no friction, the forced applied on the road by the car alone is not affected by the cars horizontal acceleration, it's just the acceleration due to gravity which we need to account for. So, the force would be Mcar(g) = 500kg(9.8m/s^2) = 4900 N